3years later amother will b 4times as old as her son,3 yrs ago,the mother's age was 2times as old as her son's agewill b 8yr hence find the age of son and mother with explainations plzz
Answer by macston(5194)

.
In 3 years, a mother will be 4 times as old as her son.
3 years ago, the mother's age was 2 times what her son's
age will be in 8 years.
.
If this is not your question, please clarify and re-post.
.
M=mother's age; S=son's age
.
Mother's age 3 years from now (M+3) is 4 times her son's
age 3 years from now (4(S+3):
.
M+3=4(S+3)
M+3=4S+12
M=4S+9
.
Mother 3 years ago (M-3) equals 2 times son 8 years from now
(2(S+8):
.
M-3=2(S+8)
M-3=2S+16
M=2S+19
.
4S+9=2S+19
2S=10
S=5
ANSWER 1: The son is 5 years old.
.
M+3=4(S+3)
M+3=4(5+3)
M+3=4(8)
M+3=32
M=29
ANSWER 2: Mother is 29 years old.
.
CHECK:
M-3=2(S+8)
29-3=2(5+8)
26=2(13)
26=26